# Embodied Red Teaming for Auditing Robotic Foundation Models

### v1

@@ Abstract
Language-conditioned robot models (i.e., robotic foundation models) enable robots to perform a wide range of tasks based on natural language instructions. Despite strong performance on existing benchmarks, evaluating the safety and effectiveness of these models is challenging due to the complexity of testing all possible language variations. Current benchmarks have two key limitations: they rely on a limited set of human-generated instructions, missing many challenging cases, and they focus only on task performance without assessing safety, such as avoiding damage. To address these gaps, we introduce Embodied Red Teaming (ERT), a new evaluation method that generates diverse and challenging instructions to test these models. ERT uses automated red teaming techniques with Vision Language Models (VLMs) to create contextually grounded, difficult instructions. Experimental results show that state-of-the-art models frequently fail or behave unsafely on ERT tests, underscoring the shortcomings of current benchmarks in evaluating real-world performance and safety. Code and videos are available at: https://sites.google.com/view/embodiedredteam.
1 Introduction
@@
Commanding robots through natural language has been a long-standing goal in robotics. Recent advances in foundational robot models (Ke et al., 2024; Wu et al., 2023; Chen et al., 2023; Brohan et al., 2022; 2023; Parakh et al., 2024; Ajay et al., 2024) trained on large-scale internet text have shown promising results in enabling robots to follow language instructions and generalize to ones not seen during training. However, despite near-perfect performance on existing benchmarks (Mees et al., 2022; James et al., 2020), these benchmarks often lack diversity in instruction phrasing and miss many commands real users might give. For example, we found that a model successfully executes "Close the drawer" but fails with "Grip the handle gently and pull it," even though both describe the same task (Section 4.1). Therefore, evaluation results on the existing benchmark may overestimate the models’ true performance in real-world scenarios where user instructions may differ from the existing benchmarking instructions.
To effectively deploy foundational robot models in the real world, we need robust benchmarking methods that capture the diverse instructions real users might give. We propose using red teaming—testing the robot model with instructions specifically designed to challenge its capabilities. These instructions are crafted to be both natural and representative of the plausible ways users might express their intentions. Thus, a model that truly generalizes should maintain high performance on these challenging instructions, whereas a drop in success rate would signal over-fitting to the instructions in the existing benchmark.
How can we effectively red team language-conditioned robot models to uncover instructions that challenge their capabilities? First, the instructions must be feasible for the robot in its real-world environment. For instance, asking a stationary manipulator to fly is unreasonable. Second, the instructions should expose the model’s weaknesses by leading it to fail at completing valid tasks. While human annotators could generate such instructions, this approach is costly and difficult to scale.
A promising alternative is leveraging large language models (LLMs) (Stiennon et al., 2020) for red teaming, as they have been effective in generating challenging instructions for chatbot models (Yu et al., 2023; Hong et al., 2024; Perez et al., 2022). However, applying these methods directly to robots is problematic since instructions must be grounded in the robot’s environment. For example, testing a robot’s ability to move a table is meaningless if no table is present. To overcome this, we propose Embodied Red Teaming (ERT): using vision-language models (VLMs) to incorporate environmental context. By reasoning over both visual and textual inputs, VLMs can generate feasible, context-aware instructions that reveal weaknesses in robot models.
Our main contribution is an automated method for auditing language-conditioned robot foundation models using VLMs. We demonstrate that ERT generates instructions that effectively challenge state-of-the-art robot models, indicating that current benchmarks overlook many instructions these robot models fail to execute. Additionally, we found that these models can exhibit unsafe behaviors, such as causing collisions or dropping objects, posing safety risks. To our knowledge, this is the first study to using red teaming to test language-conditioned robot models and identify risks from unexpected instructions.
2 Preliminaries
@@
We define the robot policy being audited as $\pi$ that takes high-dimensional sensory observations (e.g., images) $\bm{o}$ and natural language instructions $\bm{c}$ as inputs and outputs actions $\bm{a}=\pi(\bm{o},\bm{c})$ to accomplish the task described by $\bm{c}$. The common paradigm for training $\pi$ is to fine-tune pre-trained vision and language models for a set of sensorimotor tasks. As robots’ policies are derived from vision and language models pre-trained on large amounts of data, it is expected that the robots’ policies will generalize to a wide range of natural language instructions, visual scenarios, and even new tasks beyond the training set. For instance, a language-conditioned robot model should successfully execute the command "Shut off the LED light," even if it was only trained on instructions like "Turn off the light."
3 Embodied Red Teaming
Goal: The goal of embodied red teaming (ERT) for robot policies is finding a diverse set of test instructions $\{\bm{c}_{1},\cdots,\bm{c}_{N}\}$ causing the robot policy to fail, where $N$ is the instruction set size. These instructions must represent feasible tasks in the current environment and cause the robot to fail.
@@ 3 Embodied Red Teaming
Why must the instructions be diverse and feasible? Diversity ensures comprehensive auditing of robot’s performance, avoiding gaps that a narrow set of tests might leave. Feasibility ensures that the robot is being tested on tasks it is expected to perform, excluding unrealistic or irrelevant instructions.
@@
where $R(\pi,\bm{c}_{i})$ measures the robot’s performance on instruction $\bm{c}_{i}$, and Div quantifies the diversity of the selected instructions. FeasibleSet defines the conditions that make an instruction feasible, which are detailed below.
3.1 Implementation: Iterative In-Context Refinement
To optimize the objective in Equation 1, we implement ERT using vision-language models (VLMs), since the FeasibleSet includes both images and text. We refer to the VLM used for generating instructions in ERT as the “red-team VLM.” In FeasibleSet, we provide an image of the environment—indicating the entities and objects available to the robot—and a task description in language to specify the tasks for which ERT should generate instructions. Note that we did not explicitly describe the objects and entities to the red-team VLM but just provided the image of the environment.
Overview: Since VLMs lack prior knowledge about which instructions might lead the robot policy to fail, ERT refines instructions iteratively based on the robot policy’s execution results on the generated instructions. In each iteration, the red-team VLM generates $N$ instructions. ERT utilizes the VLM’s in-context learning capability to increase the likelihood of producing instructions that cause the robot to fail during successive refinement steps. In-context learning allows VLMs to adapt their output toward specific goals by including relevant examples in the prompt. In this case, instructions that previously caused robot failures are included in the prompt, guiding the red-team VLM to generate more failure-inducing instructions. To further enhance diversity, each iteration uses best-of-$M$ sampling, applying a diversity metric to select the most diverse set of instructions. Implementation details are provided below and in Algorithm 1.
@@ 3.1 Implementation: Iterative In-Context Refinement
1:Input: Metric $R$, target robot policy $\pi$, feasible set description FeasibleSet, number of refinement $K$, and budget per refinement $N$
@@
4:Initialize the prompt $\bm{x}\leftarrow\textsc{FeasibleSet}$ (see Section 3.1)
@@
9:     Append instructions $\bm{c}_{i}$ causing failure to EXAMPLES in the prompt $\bm{x}$ (see Section 3.1)
@@
We use GPT-4o as the red team VLM and denote it as $p$, though other pre-trained VLMs can also be used. The prompt $\bm{x}$ for $p$ is constructed using the following template:
{IMAGE} The attached image shows the robot’s environment. Generate a list of {N} instructions that are challenging the robot’s langauge comprehension capability for {TASK}, similar to the following examples {EXAMPLES}.
FeasibleSet: Here, {IMAGE} and {TASK} are parts of the FeasibleSet. {IMAGE} is an image depicting the environment, and {TASK} is a brief description of the desired task (e.g., Turn off the light). Instructions describe how users might command the robot to perform the task; for instance, "Switch off the light" and "Robot, please shut off the LED" are both feasible for "Turn off the light". The feasible set ensures that ERT generates language instructions for tasks the robot can perform in this environment.
Examples: {EXAMPLES} are instructions causing the robot to fail in previous iterations. For each instruction $\bm{c}$, failure of the robot policy $\pi$ is measured by a given metric function $R$, where $R(\pi,\bm{c})=0$ indicates failure and $R(\pi,\bm{c})=1$ indicates success. In this paper, the metric function is implemented as a program that determines task completion using the simulator’s states, but it can also be a classifier predicting task success. Instructions with $R(\pi,\bm{c})=0$ are added to the prompt $\theta$.
Best-of-$M$ sampling: To maximize the diversity of generated instructions, we sample $M$ sets of $N$ instructions from the red team VLM $p$, evaluate the diversity of each set, and select the set with the highest diversity for refinement step $k$. The index of the selected set is given by:
where $M=5$ in our experiments (Section 4).
4 Experiments
We evaluate ERT’s performance at discovering instructions that elicit failures or unsafe robotic behaviors.
Simulated environments: We evaluate ERT on the widely used CALVIN (Mees et al., 2022) and RLBench (James et al., 2020) benchmarks. The CALVIN benchmark includes 27 distinct tasks and 400 crowd-sourced natural language instructions, with approximately 15 instructions per task. RLBench consists of 18 tasks with 3 to 6 instructions per task. The robot policy is provided with a natural language instruction $\bm{c}$, observation of the environment from cameras $\bm{o}$ mounted on the gripper and above the robot. We evaluate the robot policies by rolling out trajectories using them from various initial states using, such as different object layouts. Figure 2 illustrates exemplar tasks from CALVIN and RLBench. Robot performance $R(\pi,\bm{c})$ for a given instruction $\bm{c}$ is measured by the success rate based on the task-specific criteria defined by the benchmark. For example, in the "close the drawer" task, the robot is instructed with a prompt such as "Hold the drawer handle and shut it". Success is determined by whether the drawer is closed.
@@ 4 Experiments
Language-conditioned robot policies: We tested the ERT-generated instructions on two state-of-the-art language-conditioned robot policies, GR-1 Wu et al. (2023) and 3D-Diffuser Ke et al. (2024), using their publicly available pre-trained checkpoints. Both models are fine-tuned on expert demonstrations. GR-1 is a transformer model pre-trained on large-scale video dataset with language instructions. Note that we did not evaluate GR-1 on RLBench, as models for RLBench were not released. 3D-Diffuser is a diffusion policy model that takes images and language embeddings as inputs, with instructions processed by pre-trained CLIP models (Radford et al., 2021). Both GR-1 and 3D-Diffuser achieved high success rates (around 90%) on the CALVIN benchmark.
4.1 Language-conditioned robots struggle with instruction generalization
Setup: We evaluate the performance of GR-1 and 3D-Diffuser on several instruction sets: those from the CALVIN benchmark and RLBench environments (denoted as Training), rephrased versions of these instructions (denoted as Rephrase), and instructions generated by ERT. The rephrased instructions are generated by providing the original benchmark instructions as example inputs to the red team VLM. We consider rephrased instructions as a naive method for generating new instructions beyond the existing benchmarks, serving as a baseline to highlight ERT’s effectiveness in discovering instructions that cause robot failures.
In our experiments, we make the instructions being tested similar to the number of instructions in the original training instruction set for each robot policy to ensure a fair comparison on the performance (success rate) of the robot’s policy. For the CALVIN environment, we prompt ERT to generate 10 instructions for each of 27 tasks, totaling 270 instructions. For 3D-Diffuser, we perform iterative refinement, as described in Algorithm 1. For GR-1, we instead re-use the instructions found from performing ERT on 3D-Diffuser for testing GR-1, aiming to see if instructions causing one model to fail also lead to failure in the other. For RLBench, we generate 3 to 6 instructions for each of 18 tasks using ERT. All generated instructions are provided in Appendix A.4 for CALVIN and Appendix A.5 for RLBench.
Metric: The models are evaluated for each instruction across all initial states defined by the benchmarks. The initial states vary across the initial configuration of the robot and the initial positions of objects in the scene. The success rate of a robot policy $\pi$ on instruction $\bm{c}$ is denoted as $R(\pi,\bm{c})$. The model’s overall performance on an instruction set $\{\bm{c}_{1},\cdots,\bm{c}_{N}\}$ is the average success rate across all instructions:
@@ 4.1 Language-conditioned robots struggle with instruction generalization
For ERT and Rephrase, we report the robots’ mean performance on instructions generated using five different random seeds, along with the 95% confidence intervals estimated via bootstrapping. Note that we cannot compute confidence intervals for the robots’ mean performance on the existing benchmark instructions because they consist of only one set of instructions. We summarize our quantitative and qualitative results below.
Failure to generalize. Figures 3(a) and 3(b) demonstrate that while GR-1 and 3D-Diffuser perform nearly optimally on the training instruction sets from the CALVIN and RLBench benchmarks, their performance degrades significantly on instructions generated by ERT. Specifically, in RLBench (Figure 3(b)), 3D-Diffuser completely fails on instructions outside its training dataset. Comparison of instructions from CALVIN and RLBench with those generated by ERT shown in Table 1 shows that ERT produces valid instructions while the robots fail to execute. These findings suggest that state-of-the-art language-conditioned robot policies overfit to the narrow set of training instructions in existing benchmarks, raising concerns about their deployment in real-world scenarios where users may phrase instructions differently.
@@
Rephrasing also degrades performance. Figure 3(a) shows that simply rephrasing existing benchmark instructions leads to a degradation in the robots’ performance, evidenced by lower success rates on rephrased instructions. GR-1 performs better than 3D-Diffuser on these rephrased instructions, suggesting that 3D-Diffuser is less capable of generalizing to rephrased instructions—possibly because GR-1 is trained on larger-scale video and language datasets.
ERT identifies more challenging instructions. As shown in Figure 3(a), rephrased instructions didn’t cause GR-1 to fail significantly caompred with the training instructions, but both 3D-Diffuser and GR-1 struggle with instructions generated by ERT—even at $k=0$ without iterative refinement—resulting in similarly low success rates. This indicates that ERT goes beyond simple rephrasing. Furthermore, as ERT’s refinement progresses (with increasing $k$) and encounters more instructions that cause the robots to fail, the generated instructions further reduce the robots’ success rates, demonstrating the effectiveness of iterative refinement.
@@
ERT generates more diverse instructions. As shown in Figure 4, ERT not only produces challenging instructions that cause the robot to fail but also generates more diverse instructions compared to other methods. Following prior works (Hong et al., 2024; Tevet & Berant, 2020), we measure instruction diversity using the BLEU score (Papineni et al., 2002) and embeddings from CLIP (Radford et al., 2021) and BERT (Reimers & Gurevych, 2019). We invert the BLEU score (1 - BLEU) and the average embedding similarity (1 - similarity) to obtain diversity scores, so higher scores indicate greater diversity. Diversity scores are calculated for instructions within the same task, and we report the average diversity in Figure 4. The BLEU diversity reflects variations in text form (e.g., wording, structure), while CLIP and BERT embedding diversities assess semantic differences. All methods exhibit similar embedding diversities, which is expected since they describe the same tasks and thus have similar semantics. However, ERT achieves significantly higher BLEU diversity, indicating that it phrases instructions in more varied ways than existing benchmarks and their rephrased versions. This suggests that ERT does not merely exploit a single vulnerability of the robot but challenges it with a broader range of instruction phrasings.
Refined instructions transfer to other models. We found that instructions causing one robot model to fail often lead to failures in other models. ERT($k=0$) instructions are generated by sampling from the VLM without examples, while ERT($k=1$) and ERT($k=2$) generate instructions conditioned on the lowest-performing instructions from the previous iteration. Although we refined ERT using only 3D-Diffuser’s success rates (see Algorithm 1), ERT($k=1$) and ERT($k=2$) also significantly reduced GR-1’s success rates. This suggests that instructions challenging for 3D-Diffuser are similarly challenging for GR-1, indicating that current state-of-the-art robot models may struggle with similar types of instructions.
Template-generated instructions could be problematic. We hypothesize that the significant performance drop of 3D diffusers on ERT-generated instructions in RLBench is due to RLBench’s template-based instructions, in contrast to CALVIN’s human-generated ones. As shown in Table 1, template-generated instructions from RLBench lack the variety of human phrasing and appear monolithic. Training 3D-Diffuser on CALVIN’s human-like instructions attains better generalization, whereas training on RLBench’s template-generated instructions significantly weakens it. This suggests that relying on templates to train language-conditioned robot policies is problematic.
@@
turn_on_lightbulb turn on the yellow light Move the robot arm to the switch and toggle it to turn off the lightbulb.
@@
Table 1: Example instructions from CALVIN, RLBench, and ERT (our method) for various tasks. ERT generates more complex yet reasonable instructions for the same tasks.
@@
Figure 4: Instruction Diversity. BLEU diversity captures variations in text form, while CLIP and BERT diversity measure semantic differences. Since all instructions describe the same task, semantic diversity is similar across methods. However, ERT achieves higher BLEU diversity, indicating it can generate more varied phrasing for the same task compared to other methods. See Section 3(a).
4.2 Large Models Struggle At Instruction Generalization

### v2

@@ Abstract
Language-conditioned robot models have the potential to enable robots to perform a wide range of tasks based on natural language instructions. However, assessing their safety and effectiveness remains challenging because it is difficult to test all the different ways a single task can be phrased. Current benchmarks have two key limitations: they rely on a limited set of human-generated instructions, missing many challenging cases, and focus only on task performance without assessing safety, such as avoiding damage. To address these gaps, we introduce Embodied Red Teaming (ERT), a new evaluation method that generates diverse and challenging instructions to test these models. ERT uses automated red teaming techniques with Vision Language Models (VLMs) to create contextually grounded, difficult instructions. Experimental results show that state-of-the-art language-conditioned robot models fail or behave unsafely on ERT-generated instructions, underscoring the shortcomings of current benchmarks in evaluating real-world performance and safety. Code and videos are available at: https://s-karnik.github.io/embodied-red-team-project-page/.
(a) Sensitive to instruction phrasing
(b) Overview of Embodied Red Teaming (ERT)
Figure 1: (a) Users may express intentions in various ways, but language-conditioned robots often succeed with some instructions and fail with others, highlighting their limited ability to generalize. (b) Each task can be described by multiple instructions. Embodied Red Teaming (ERT) uses a task description and environmental observations (e.g., camera images) to generate instructions that are likely to cause the robot to fail. By incorporating feedback from the robot’s execution results, ERT refines the instruction generation process to increase the probability of failure.
11footnotetext: 1 Equal contribution. Correspondence to {skarnik, zwhong}@mit.edu
I Introduction
@@
Commanding robots via natural language instructions has been a long-standing goal in robotics [22]. Recent advances in language-conditioned robot models [9, 25, 4, 2, 3, 15, 1] leveraging large-scale internet text have shown promising results in enabling robots to follow language instructions. However, despite the near-perfect performance of current models on existing benchmarks [12, 8], these models are not ready for real-world use. The main reason is that current models are sensitive to how instructions are phrased, as illustrated in Figure 1(a), an issue existing evaluations overlook. For instance, when we command a robot to open the drawer, a model successfully executes "Open the drawer" but fails to execute when the instruction is rephrased as "Grip the handle gently and pull it towards yourself," (Section IV-A). As a result, strong performance on language instructions on current benchmarks does not ensure satisfactory performance in real-world scenarios.
To effectively deploy language-conditioned robot models, we, therefore, need robust evaluation methods that capture the diverse ways in which a user may express a task to the robot in real-life interactions. The robot should function properly according to any natural language instruction that describes the task it can perform. For instance, if a robot has been trained to open drawers, we expect it to succeed on all language instructions describing the task “open drawer”. The central goal of such evaluation is to identify instructions the robot can execute but still fails to execute because of how the instruction is expressed by the user. We call the identification of such instructions as Embodied red teaming. Model developers can then include these instructions in the training set to ensure the robot model doesn’t fail because the instruction was expressed differently (i.e., apply a “patch” to the robot model). While human annotators could generate diverse instructions, this approach is costly and difficult to scale.
A promising alternative to human annotators is to use large language models (LLMs) [7, 16] to generate instructions that challenge a robot on the task it is trained to perform. LLMs have successfully been used to generate prompts that challenge chatbot models [26, 7, 16]. However, simply applying existing methods to robots is insufficient for handling the wide range of instructions users might provide. For one, LLMs cannot process images and, therefore, cannot generate instructions referring to objects in the environment. For example, the instructions pick up a coke can and “Pick up the stuff on the table” both correspond to the same task if the only object on the table is a can. Generating such context-aware instructions requires knowledge of the robot’s environment inferred through the robot’s sensory observations.
We introduce Embodied Red Teaming (ERT) whose goal is to audit robot models by creating an instruction set that (i) is grounded in the environment, (ii) is feasible for the robot to execute, (iii) but challenges the robot’s performance and (iv) is diverse enough to ensure sufficient coverage in the instruction space that provides greater confidence about the robot’s real-world performance. To achieve both (i) and (ii), ERT leverages VLMs that can analyze the environment using visual observations and in the context of the task description provided in natural language. Because VLMs are not trained to exploit the robot’s weaknesses, they are ineffective at generating challenging instructions out-of-the-box. To create more challenging instructions, ERT iteratively refines an instruction set using in-context learning [13, 5] and feedback from the robot’s execution. To ensure that the instructions are diverse, we use the VLM to sample multiple sets of instructions and select the set with the highest diversity.
Our contribution is in exposing the gap between existing benchmarks and real-world use cases for language-conditioned robot models. Current benchmarks rely on instructions curated by human annotators to evaluate these models. However, we demonstrate that ERT generates instructions for tasks the robot can perform, but the state-of-the-art language-conditioned robot models fail to complete. This reveals that existing evaluation benchmarks do not reflect robot’s real world performance. To the best of our knowledge, this is the first study to use red teaming to test language-conditioned robot models.
II Preliminaries
@@
Let the robot policy being audited be $\pi$ that takes high-dimensional sensory observations (e.g., images) $\bm{o}$ and natural language instructions $\bm{c}$ as inputs and outputs actions $\bm{a}=\pi(\bm{o},\bm{c})$ to accomplish the task described by $\bm{c}$. The common paradigm for training $\pi$ leverages pre-trained vision and language models, and therefore, it is expected that the robots’ policies will generalize to a wide range of natural language instructions, visual scenarios, and even new tasks beyond the training set. For instance, a language-conditioned robot model should successfully execute the command “Shut off the LED light,” even if it was only trained on instructions like “Turn off the light.”
III Embodied Red Teaming
Goal: The goal of embodied red teaming (ERT) is to find a diverse set of feasible instructions $\{\bm{c}_{1},\cdots,\bm{c}_{N}\}$ which the robot fails to execute, where $N$ is the instruction set size. These instructions must describe feasible tasks that are both (i) achievable within the current environment (e.g., instructions involving pushing a table require a table to be present) and (ii) within the scope of tasks the robot has been trained to perform.
@@ III Embodied Red Teaming
Why must the instructions be diverse and feasible? Diversity ensures comprehensive auditing of robot’s performance, avoiding gaps that a narrow test suite might leave. Feasibility ensures that the robot is being tested on tasks it is trained to perform in the current environment.
@@
where $R(\pi,\bm{c}_{i})$ measures the robot’s performance on instruction $\bm{c}_{i}$, and Div quantifies the diversity of the selected instructions. FeasibleSet defines the conditions that make an instruction feasible and is detailed below.
III-A Implementation: Iterative In-Context Refinement
To optimize the objective in Equation 1, we implement ERT using vision-language models (VLMs) since we rely on both images and text to describe the FeasibleSet that indicate which instructions the robot can perform in the current environment. Algorithm 1 outlines the implementation of ERT. We refer to the VLM used for generating instructions in ERT as the “red-team VLM.”. In this paper, we use GPT-4o as the red-team VLM and denote it as $p$, though other pre-trained VLMs can also be used. The prompt $\bm{x}$ for $p$ is constructed using the following template:
{IMAGE} The attached image shows the robot’s environment. Generate a list of {N} instructions that are challenging the robot’s language comprehension capability for {TASK}, similar to the following examples {EXAMPLES}.
FeasibleSet: Here, {IMAGE} and {TASK} define the FeasibleSet. {IMAGE} is an image depicting the environment, and {TASK} is a brief and under-specified description of the desired task (e.g., turn_off_light). Instructions describe how users might command the robot to perform the task; for instance, "Switch off the light" and "Robot, please shut off the LED" are both feasible user instructions for the task turn_off_light. Prompting $p$ to generate feasible instructions ensures that these instructions are limited to the tasks the robot can perform in the given environment. Note that we do not explicitly describe the objects and their relationships to the red-team VLM but provide an image of the environment.
EXAMPLES: VLMs lack prior knowledge of which instructions may cause the robot models to fail. To address this, ERT iteratively refines instructions based on the robot’s execution results. In each iteration, the red-team VLM generates $N$ instructions and selects those that lead to failures as the {EXAMPLES} in the prompt. Since VLMs are capable of doing in-context learning [5], including failure examples increases the likelihood of sampling similar but different instructions at the next generation due to the stochastic nature of VLMs. For each instruction $\bm{c}$, the robot model $\pi$ is evaluated using a reward function $R$, where $R(\pi,\bm{c})=0$ indicates failure and $R(\pi,\bm{c})=1$ indicates success. In this study, $R$ is implemented as a program that assesses task completion based on the simulator’s states, but it can also be a classifier that predicts task success. Instructions with $R(\pi,\bm{c})=0$ are added to the prompt $\theta$.
Best-of-$M$ sampling: To maximize the diversity of generated instructions, we sample $M$ sets of $N$ instructions from the red-team VLM $p$, evaluate the diversity of each set, and select the set with the highest diversity for refinement step $k$. The index of the selected set is given by:
where $M=5$ in our experiments (Section IV). We define the diversity of a set of instructions as the average cosine embedding distance:
$\text{Div}(\{\bm{c}^{m}_{1},\cdots,\bm{c}^{m}_{N}\})=\dfrac{1}{2N}\sum^{N}_{i=1}\sum^{N}_{j=1}\dfrac{\phi(\bm{c}^{m}_{i})\cdot\phi(\bm{c}^{m}_{j})}{\lVert\phi(\bm{c}^{m}_{i})\rVert\lVert\phi(\bm{c}^{m}_{j})\rVert},$ (3)
where we implemented $\phi$ with CLIP embedding [17].
@@ III-A Implementation: Iterative In-Context Refinement
1:Input: Reward $R$, target robot policy $\pi$, feasible set description FeasibleSet, number of refinement $K$, and budget per refinement $N$
@@
4:Initialize the prompt $\bm{x}\leftarrow\textsc{FeasibleSet}$ (see Section III-A)
@@
9:    Append instructions $\bm{c}_{i}$ causing failure to EXAMPLES in the prompt $\bm{x}$ (see Section III-A)
@@
IV Experiments
We evaluate ERT on the widely used CALVIN [12] and RLBench [8] benchmarks. The CALVIN benchmark includes 27 distinct tasks and 400 crowd-sourced natural language instructions, with approximately 15 instructions per task. RLBench consists of 18 tasks with 3 to 6 instructions per task. The robot model is provided with a natural language instruction $\bm{c}$, observation of the environment from cameras $\bm{o}$ mounted on the gripper and above the robot. We evaluate the robot models by rolling out trajectories using them from various initial states, such as different object layouts. Figure 2 illustrates exemplar tasks from CALVIN and RLBench. Robot performance $R(\pi,\bm{c})$ for a given instruction $\bm{c}$ is measured by the success rate based on the task-specific criteria defined by the benchmark. For example, in the “close the drawer” task, the robot is instructed with a prompt such as “Hold the drawer handle and shut it.” Success is determined by whether the drawer is closed.
@@ IV Experiments
IV-A Language-conditioned robots struggle with instruction generalization
Setup: We tested the ERT-generated instructions on the state-of-the-art language-conditioned robot model 3D-Diffuser [9], using their publicly available checkpoints. The 3D-Diffuser is a diffusion policy model that takes images and language embeddings as inputs, with instructions processed by a pre-trained CLIP model [17]. The 3D-Diffuser model was trained on expert demonstrations for the tasks in CALVIN and RLBench. We evaluate the performance of 3D-Diffuser on several instruction sets below:
  • Training: Instructions from the CALVIN benchmark and RLBench environments.
  • Rephrase: These rephrased instructions were created by inputting the original benchmark instructions (Training) into an LLM (GPT-4o) as examples. Note that this is different from ERT since ERT does not take the original benchmark instructions as inputs. We use this method as a simple baseline for generating new instructions beyond existing benchmarks. For further details, see Appendix A-A.
  • ERT (Ours): We evaluated ERT using varying numbers of refinement iterations (denoted as $k$) to demonstrate its effectiveness both with and without refinement.
For evaluation, we generate new instruction sets where each is roughly equivalent in size to the original training set. In the CALVIN environment, we use ERT to generate 10 instructions for each of 27 tasks (totaling 270 instructions) and then apply three iterative refinement steps (i.e., $k\in\{1,2,3\}$, see Algorithm 1). In contrast, for RLBench, we generate 3 to 6 instructions per each of 18 tasks using ERT without refinement since the unrefined instructions already challenge the 3D-Diffuser model. All generated instructions are available in Appendix A-D for CALVIN and Appendix A-E for RLBench.
Metric: The models are evaluated for each instruction across all initial robot and object configurations as is the standard practice in the benchmarks used for evaluation. The success rate of a robot policy $\pi$ on instruction $\bm{c}$ is denoted as $R(\pi,\bm{c})$. The model’s overall performance on an instruction set $\{\bm{c}_{1},\cdots,\bm{c}_{N}\}$ is the average success rate across all instructions:
@@ IV-A Language-conditioned robots struggle with instruction generalization
We report the mean performance on the Training instructions. For ERT and Rephrase, we report the robots’ mean performance on instructions generated using five different random seeds, along with the 95% confidence intervals estimated via bootstrapping. We summarize our quantitative and qualitative results below.
@@
Robots failing to generalize: As shown in Figure 3, the 3D-Diffuser model struggles to generalize across instructions, as demonstrated by significant performance drops of the robot models on instructions generated by ERT compared to those in existing benchmarks (Training). While the 3D-Diffuser achieves high success rates on CALVIN (90%) and RLBench (79%) training instructions, its performance degrades to 53% and 3%, respectively, on ERT-generated instructions. Rephrasing existing benchmark instructions (Rephrase) also reduces success rates (66.9% vs. 92%) in the CALVIN benchmark, highlighting the sensitivity of the robot model to minor variations in phrasing. These findings reveal that current robot models overfit to narrow training instruction sets. We hypothesize that the significant performance drop of the 3D-Diffuser on ERT-generated instructions in RLBench is due to RLBench’s use of template-based instructions, unlike CALVIN’s human-generated ones. As shown in Table I, RLBench’s template-generated instructions lack the diversity of human phrasing and appear monolithic. Training the 3D-Diffuser on CALVIN’s human-like instructions improves generalization while training on RLBench’s template-based instructions hinders it.
ERT v.s. Rephrase: Figure 3(a) shows that ERT, even without refinement, produces more challenging instructions than Rephrase, leading to lower success rates. This indicates that giving the red-team model an under-specified task description (e.g., open_drawer) is more effective for generating challenging instructions than merely rephrasing existing ones. We hypothesize that including the original instructions in the LLM prompt causes the generated instructions to resemble the training data for the robot models. Additionally, iterative refinement further lowers the robot model’s success rates, demonstrating its effectiveness in exploiting the models’ weaknesses.
@@
ERT generates more diverse instructions: As shown in Figure 4, ERT not only produces challenging instructions that cause the robot to fail but also generates more diverse instructions compared to other methods. Following prior works [7, 23], we measure instruction diversity using the BLEU score [14] and embeddings from CLIP [17] and BERT [18]. We invert the BLEU score (1 - BLEU) and the average embedding similarity (1 - similarity) to obtain diversity scores, so higher scores indicate greater diversity. Diversity scores are calculated for instructions within the same task, and we report the average diversity in Figure 4. The BLEU diversity reflects variations in text form (e.g., wording, structure), while CLIP and BERT embedding diversities assess semantic differences. All methods exhibit similar embedding diversities, which is expected since they describe the same tasks and thus have similar semantics. However, ERT achieves significantly higher BLEU diversity, indicating that it phrases instructions in more varied ways than existing benchmarks and their rephrased versions. This suggests that ERT does not merely exploit a single instruction causing the robot to fail but challenges the robot models with a broader range of instruction phrasings.
@@
turn_on_lightbulb turn on the yellow light Move the robot arm to the switch and toggle it to turn on the lightbulb.
@@
TABLE I: Example instructions from CALVIN, RLBench, and ERT (our method) for various tasks. ERT generates more complex yet reasonable instructions for the same tasks. Note that not every instruction causes the robot to fail.
@@
Figure 4: Instruction Diversity. BLEU diversity captures variations in text form, while CLIP and BERT diversity measure semantic differences. Since all instructions describe the same task, semantic diversity is similar across methods. Regardless of the number of refinement $k$, ERT achieves higher BLEU diversity, indicating it generates more varied phrasing for the same task compared to other methods. See Section 3(a).
IV-B Failed Instructions transfer across robot models
Setup: We examined whether instructions that cause one robot model (3D-Diffuser) to fail similarly cause failure of another state-of-the-art model, GR-1—a transformer-based model pre-trained on a large-scale video dataset with language instructions. (Note: GR-1 was not evaluated on RLBench due to unavailable benchmarks.)
Results: As Figure 3(a) shows, instructions leading to failures in the 3D-Diffuser also frequently caused failures in GR-1. Although the ERT refine the instructions based solely on the 3D Diffuser’s performance (see Algorithm 1), both ERT($k=1$) and ERT($k=2$) with refinement significantly reduced GR-1’s success rates as well. This suggests that instructions challenging for the 3D Diffuser also challenge GR-1, highlighting a common vulnerability among state-of-the-art robot models. Notably, GR-1 performed better on rephrased instructions, likely due to its training on a more diverse video and language dataset, whereas the 3D Diffuser was limited to the CALVIN and RLBench datasets.
IV-C Large models also struggle at instruction generalization